Find the three arithmetic means between 21 and 13.
Answer by venugopalramana(3286)

Find the three arithmetic means between 21 and 13.
LET THE MEANS BE
21,P,Q,R,13
T5=13=21+(5-1)D=21+4D
4D=13-21=-8
D=-2
P=21-2=19
Q=19-2=17
R=17-2=15